Purdue University MA265 Linear Algebra Spring 2012 Syllabus Instructor: Uli Walther Office: 746 MATH Phone: 49-41959

                                E-mail: walther@math.purdue.edu

1 General Information

1.1 Class time / location:

TTh 9:00 -10:15 REC 226.

1.2 Office Hours:

Tue 12:00-1:00 and Th 1:00-2:00.

1.3 Textbook:

Elementary Linear Algebra B. Kolman, D. Hill 9th Edition, Prentice Hall 2007.

Lab Manual: Linear Algebra with Matlab D. Hill, D. Zitarelli 3rd edition, Prentice Hall 2004.

Matlab files and Supplementary exercises: see department course web page

http://www.math.purdue.edu/academic/courses/MA26500/.

1.4 Webpages:

The department course webpage is http://www.math.purdue.edu/ma26500.

This page contains matlab files to download for you. These will be needed during the course for homework.

Our class page is http://www.math.purdue.edu/~walther/teach. It includes

syllabus, some other info concerning the class, and a link to our Project Rhea (see below).

The WebAssign student login page is http://www.webassign.net/purdue/login.html.

1.5 Statement for Students with Disabilities

See the webpage, http://www.math.purdue.edu/~walther/teach.

1.6 Calculators

Calculators are not allowed on any test.

2 Course Structure

If you disagree with the grading in any category, you need to resubmit the item in question with a written explanation stating why (and not just that) you deserve more credit.

2.1 Homework:

Most HW is submitted online by Thursday evening 11:40pm. The few problems not submitted online, including all Matlab problems, will be collected every Thursday in class or in my office no later than 2:00pm. No late assignments will be accepted. Neither will homework deposited anywhere else.

The homework due in any given week is the homework corresponding to the material of the previous week. Homework must be readable and must be stapled. Illegible scribblings will receive no credit from the grader.

You are encouraged to attempt all the questions and discuss with your classmates. However, the write-up must be of your own.

2.2 Project Rhea

You can receive up to 5% extra credit from your our contributions to Rhea. For ideas what to contribute, see the Rhea page. This is due on the day of the final.

2.3 Midterms:

There will be two midterms, Thursday, Feb 16 and Thursday, Mar 29. Makeup tests may be given in extraordinary instances, but only with documented reasons.

2.4 Final Exam:

The Final Exam will be given during the Final Exam Week.

The location and date of the final will be announced in class.

All exams will be multiple choice.

2.5 Course Grade:

Your course grade will be determined using the following distributions:

Rhea  up to 5%
HW  25%
2 Midterms, each 20%
Final  35%